OpenStudy (anonymous):

1. Let f(x)=3x-10. Find the value of f(-3) a. 21 b. -1 c. -19 2. Find the slope of the line that passes through (3,-4) & (-1,2). a. 2/3 b. 3/2 c. -2/3 d. -3/2 3. Find the slope of the line passing through (4,9) & (-2,9). a. 0 b. underfined

OpenStudy (anonymous):

Please, please help.

OpenStudy (anonymous):

\[f(x)=3x-10\] \[f(-3)=3\times (-3)-10\]

OpenStudy (anonymous):

And, do I just solve it like that?

OpenStudy (anonymous):

Find the slope of the line passing through (4,9) & (-2,9). since the \(y\) values are the same, the slope is \(0\)

OpenStudy (anonymous):

yes, compute \(3\times (-3)-10\) and that will be your answer

OpenStudy (anonymous):

Oh, okay, thank you! So, so much! Can you do the third one?

OpenStudy (anonymous):

Sorry, 2nd

OpenStudy (anonymous):

i wrote the answer to the third one above

OpenStudy (anonymous):

\((3,-4) , (-1,2)\) slope is \[m=\frac{-4-2}{3-(-1)}\] compute that number

OpenStudy (anonymous):

Thank you, thank you, thank you!

OpenStudy (anonymous):

yw btw you do it carefully, you should get \(\frac{-6}{4}=-\frac{3}{2}\)

OpenStudy (anonymous):

Oh, okay! I forgot to simplfy.

OpenStudy (anonymous):

The x in 3x(-3)-10 I'm confused what I do with it